The blue curve on the following graph represents the demand curve facing a firm that can set its own prices Use the graph input tool to help you answer the following questions. You will not be graded on any changes you make to this graph Note: Once you enter a value in a white field, the graph and any corresponding amounts in each grey field will change accordingly Graph Input Tool Market for Goods Quantit 25 Demanded (Units) Demand Price (Dollars per unit) 50.00 a 60 50 and 0 5 10 15 20 25 30 35 40 45 50 QUANTITY (Units)
Answer #1
Output (Q) P($) TR($) = P*Q MR($)=∆TR/∆Q
0 100 0
10 80 800 80
20 60 1200 40
25 50 1250 10
30 40 1200 -10
40 20 800 -40
50 0 0 -80

1375 1250 1125 1000 875 750 625 TR 500 375 250 125 0 40 0 5 15 20 25 30 35 45 50 55 Quantity (Number of units) 10 LO Total Re

TR of producing 9 units = $82 * 9 = $738

TR of producing 10 units = $80 * 10 = $800

The Marginal revenue of 10th unit produced = 800 - 738 = $62

TR of producing 19 units = $62 * 19 = $1178

TR of producing 20 units = $60 * 20 = $1200

The Marginal revenue of 20th unit produced = 1200 - 1178 = $22

100 80 60 40 20 MR 30 45 5 10 15 20 25 40 -20 -40 -60 Quantity (Units) Marginal Revenue($) 35

Answer to the blank: decreasing